# Proactive Model Adaptation Against Concept Drift for Online Time Series Forecasting

### v4

@@ 4.5. Discussion
Prior approaches to online time series forecasting perform model adaptation based on feedback in forecasting previous samples, which mainly focus on learning recent data patterns. In Table 2, we compare existing methods and Proceed in terms of the training data and the model adaptation techniques at each online time $t$. Most methods use one newly acquired training sample, while SOLID (Chen et al., 2024) selects more recent samples that share similar lookback windows with the test sample and are assumed to share a similar concept. SOLID and Proceed simply fine-tune the forecast model by gradient descent, while FSNet and OneNet further generate additional parameter adjustments based on the forecasting feedback. Since the feedback is delayed by $H$ steps and cannot reveal the test concept, we propose a novel step called proactive model adaptation which aims to mitigate the effects of concept drift between the training sample and the test sample. As this core step is orthogonal to existing methods, Proceed can incorporate the data augmentation technique of SOLID and the feedback-based model adaptation techniques of FSNet and OneNet.
@@ 5.1. Experimental Settings
Forecast Models. As our framework is model-agnostic, we pretrain three popular and advanced forecasting models, including TCN (Bai et al., 2018), PatchTST (Nie et al., 2023), and iTransformer (Liu et al., 2024). We report the forecasting errors of these pretrained models without any online learning method (Method=” \” in Table 3-4).
@@ 5.2.1. Effectiveness
To verify the effectiveness of our proposed framework, we compare the predictive performance of Proceed and other baselines on the five popular datasets. We repeat each experiment with 3 runs and report the average results. As shown in Table 3-4, Proceed achieves the best performance in most cases, reducing the average forecast errors of all models without online learning by 42.3%, 10.3%, 12.9% for TCN, PatchTST, and iTransformer, respectively. Also, it is worth mentioning that all forecast models have been enhanced by RevIN (Kim et al., 2022), a data adaptation approach to concept drift by reducing the distribution shifts w.r.t. the mean and standard deviation of time series. Since all online learning methods can still achieve remarkable improvements, it suffices to support our claim that time series forecast models need model adaptation to handle more complex concept drifts.
@@ Appendix A Theoretical Analysis
Inspired by a recent work (Bai et al., 2023), we can prove that it is feasible for the forecast model with proactive model adaptation to have lower forecasting error than one without proactive model adaptation.
@@ B.1. Online Model Adaptation
Online model adaptation, or online learning (Hoi et al., 2021), has been a popular learning paradigm that updates models on new data instantly or periodically. In the general field, most efforts focus on addressing the catastrophic forgetting issue (McCloskey and Cohen, 1989) stemming from excessive updates, and numerous continual learning (Wang et al., 2023) methods have been developed to retain acquired knowledge of past data by rehearsal mechanisms (Buzzega et al., 2020) and regularization terms (Schwarz et al., 2018), which can be seamlessly incorporated into our framework. Most recently, SOLID (Chen et al., 2024) proposed to adapt the forecast model by several selected training samples which are assumed to have a similar context (i.e., concept in this paper) to each test sample. SOLID (Chen et al., 2024) relies on heuristic measures about context similarity and is not fully aware of unobserved contexts. The selected training samples may not share the exactly same concept with the test sample. Thus it is desirable to further adopt our proposed framework. We would like to leave the combination of training data sampling and proactive model adaptation as future work.
After the submission of our manuscript, there are two concurrent researches (yee Ava Lau et al., 2025; Liang et al., 2024) that also noticed the information leakage in FSNet and OneNet. Since the ground-truth $H$-step values of the last prediction are not fully observed, both works propose to generate pseudo labels of the unobserved values, concatenate the pseudo labels with observed labels to simulate complete ground truth, and calculate the forecast errors to update the model. The potential drawback is that the pseudo label generator is still susceptible to concept drift and may produce low-quality pseudo labels, leaving the concept drift issue unresolved.
On top of online learning, rolling retraining (Li et al., 2022) is another learning paradigm that periodically re-trains a new model from scratch on all historical data, while the cost of frequent retraining is unaffordable. In practice, we can perform rolling retraining once a month and adopt daily updates during the interval to learn new patterns timely. Our work is orthogonal to rolling retraining.
@@ B.2. Data Adaptation
Apart from model adaptation, data adaptation is another mainstream approach to concept drift in time series forecasting, which is orthogonal to model adaptation. The goal of data adaptation is to normalize historical training data and future test data into a common data distribution (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023), reducing overfitting risks. Among them, RevIN (Kim et al., 2022) is the most popular method that applies instance normalization to each time series lookback window and restores the statistics to the corresponding predictions, making each sample follow a similar distribution. Such normalization-based data adaptation methods mainly focus on the statistical changes in the mean and deviation of time series, while they overlook the distribution shifts in more complex temporal dependencies between time steps and spatial dependencies between variates (Zhao and Shen, 2024). Meanwhile, the forecast model may underfit the normalized time series, as the removed statistics can serve as informative signals for prediction (Liu et al., 2022).
@@ C.1. Datasets
  • ETT (Zhou et al., 2021) (Electricity Transformer Temperature) records seven features of electricity transformers from 2016 to 2018. ETTh1 is hourly collected and ETTm1 is 15-minutely collected.
  • ECL (Zeng et al., 2023) records the hourly electricity consumption of 321 clients from 2012 to 2014.
  • Weather (Zhou et al., 2021) records 21 features of weather, e.g., air temperature and humidity in 2020.
  • Traffic (Wu et al., 2020) records the hourly road occupancy rates recorded by the sensors of San Francisco freeways from 2015 to 2016.
@@ C.2. Evaluation Settings
To the best of our knowledge, there is no open-source evaluation framework that considers the temporal gap in online learning for time series forecasting. Though Zhang et al. (Zhang et al., 2023) noticed the feedback delay, they sidestep this issue by performing one prediction task every $H$ time steps and performing the next task until the ground truth is available. Concretely, they use $\mathbf{X}_{t-L:t}$ to forecast $\mathbf{X}_{t+1:t+H}$ and then use $\mathbf{X}_{t-L+H:t+H}$ to forecast $\mathbf{X}_{t+H+1:t+2H}$. In contrast, the common practice is to perform forecasting at each time step between $t$ and $t+H$, continually updating the previous predictions of $\mathbf{X}_{t-L+H:t+H}$. The reason is that short-term forecasting is typically easier than long-term forecasting. Therefore, in this work, we implement a more realistic evaluation framework for online learning where we perform forecasting at each time step, which remains consistent with the traditional evaluation setting.
As for the ETTm1 dataset, popular works (e.g., Informer (Zhou et al., 2021), PatchTST (Nie et al., 2023), and iTransformer (Liu et al., 2024)) usually use 69,680 time steps of the dataset. In contrast, FSNet and OneNet only used a quarter of it, i.e., 17,420 steps, and the training set only contains 3484 samples. This setting is in favor of small models (e.g., TCN) but is challenging for the state-of-the-art time series forecast models. Therefore, we take all 69,680 time steps following PatchTST and iTransformer. As such, the ETTm1 dataset in our experiment has the greatest number of samples, enriching the diversity of our datasets.
@@ C.3. Implementation Details
We conduct experiments on four Nvidia 4090 24GB GPUs. We ran each experiment 3 times with different random seeds and reported the average results. We follow the official implementation of the forecast models and the online learning methods, using their recommended model hyperparameters (e.g., the number of layers) and the Adam optimizer. In cases where default hyperparameter values are not provided, we conduct a grid search to find the optimal hyperparameters that yield the best performance. Following FSNet and OneNet, the lookback length $L$ for TCN is set to be $60$. For PatchTST and iTransformer, we search for the optimal lookback length and set $L$ to be $512$. We enhance TCN and FSNet by RevIN (Kim et al., 2022) which is widely adopted by the state-of-the-art forecast models to reduce distribution shifts, lowering the forecasting loss in our experiments. As for the Optimal and Practical variants, we report the lowest errors when applying or not applying RevIN to FSNet and OneNet. Empirically, the Optimal variant without RevIN achieves better performance in some cases.
@@ E.2. Performance with Partial Ground Truth
In this section, we introduce variants that update the model with feedback on $\mathbf{Y}_{t-H}$ and partial ground truth $\{\tilde{\mathbf{Y}}_{t-i}\}_{i=1}^{H-1}$, where ${\tilde{\mathbf{Y}}_{t-i}=[\mathbf{v}_{t-i},\cdots,\mathbf{v}_{t-1}]}$. Since each update involves $H$ training samples, the GPU memory overhead becomes about $H$ times than one involving $\mathbf{Y}_{t-H}$ only. We leave out results on ECL and Traffic datasets, where the GPU memory overhead exceeds the limit of our Nvidia 4090 GPUs (24GB). As shown in Table 9, Proceed consistently outperforms the online gradient descent method under this setting. Compared with Table 3, using both $\mathbf{Y}_{t-H}$ and partial ground truth for feedback-based adaptation can improve the performance, while the improvement is insignificant. This indicates that samples with partial ground truth cannot reveal the concept of the test sample, necessitating proactive model adaptation. Though the partial ground truth can be beneficial to our method, the finetuning cost scales up with $H$ (e.g., up to either 96$\times$ latency or 96$\times$ GPU memory). Thus, it would be much more efficient if using $\mathbf{Y}_{t-H}$ only. Notably, our proposed method using $\mathbf{Y}_{t-H}$ only can still outperform GD using $H$ training samples in most cases. .

### v5

@@ 4.5. Discussion
Prior approaches to online time series forecasting perform model adaptation based on feedback in forecasting previous samples, which mainly focus on learning recent data patterns. In Table 2, we compare existing methods and Proceed in terms of the training data and the model adaptation techniques at each online time $t$. Most methods use one newly acquired training sample, while SOLID (SOLID) selects more recent samples that share similar lookback windows with the test sample and are assumed to share a similar concept. SOLID and Proceed simply fine-tune the forecast model by gradient descent, while FSNet and OneNet further generate additional parameter adjustments based on the forecasting feedback. Since the feedback is delayed by $H$ steps and cannot reveal the test concept, we propose a novel step called proactive model adaptation which aims to mitigate the effects of concept drift between the training sample and the test sample. As this core step is orthogonal to existing methods, Proceed can incorporate the data augmentation technique of SOLID and the feedback-based model adaptation techniques of FSNet and OneNet.
@@ 5.1. Experimental Settings
Forecast Models. As our framework is model-agnostic, we pretrain three popular and advanced forecasting models, including TCN (TCN), PatchTST (PatchTST), and iTransformer (iTransformer). We report the forecasting errors of these pretrained models without any online learning method (Method=” \” in Table 3-4).
@@ 5.2.1. Effectiveness
To verify the effectiveness of our proposed framework, we compare the predictive performance of Proceed and other baselines on the five popular datasets. We repeat each experiment with 3 runs and report the average results. As shown in Table 3-4, Proceed achieves the best performance in most cases, reducing the average forecast errors of all models without online learning by 42.3%, 10.3%, 12.9% for TCN, PatchTST, and iTransformer, respectively. Also, it is worth mentioning that all forecast models have been enhanced by RevIN (RevIN), a data adaptation approach to concept drift by reducing the distribution shifts w.r.t. the mean and standard deviation of time series. Since all online learning methods can still achieve remarkable improvements, it suffices to support our claim that time series forecast models need model adaptation to handle more complex concept drifts.
@@ Appendix A Theoretical Analysis
Inspired by a recent work (DRAIN), we can prove that it is feasible for the forecast model with proactive model adaptation to have lower forecasting error than one without proactive model adaptation.
@@ B.1. Online Model Adaptation
Online model adaptation, or online learning (Hoi2021), has been a popular learning paradigm that updates models on new data instantly or periodically. In the general field, most efforts focus on addressing the catastrophic forgetting issue (Catastrophic) stemming from excessive updates, and numerous continual learning (Wang2023) methods have been developed to retain acquired knowledge of past data by rehearsal mechanisms (DER) and regularization terms (EWC), which can be seamlessly incorporated into our framework. Most recently, SOLID (SOLID) proposed to adapt the forecast model by several selected training samples which are assumed to have a similar context (i.e., concept in this paper) to each test sample. SOLID (SOLID) relies on heuristic measures about context similarity and is not fully aware of unobserved contexts. The selected training samples may not share the exactly same concept with the test sample. Thus it is desirable to further adopt our proposed framework. We would like to leave the combination of training data sampling and proactive model adaptation as future work.
After the submission of our manuscript, there are two concurrent researches (DSOF; ActNow) that also noticed the information leakage in FSNet and OneNet. Since the ground-truth $H$-step values of the last prediction are not fully observed, both works propose to generate pseudo labels of the unobserved values, concatenate the pseudo labels with observed labels to simulate complete ground truth, and calculate the forecast errors to update the model. The potential drawback is that the pseudo label generator is still susceptible to concept drift and may produce low-quality pseudo labels, leaving the concept drift issue unresolved.
On top of online learning, rolling retraining (DDGDA) is another learning paradigm that periodically re-trains a new model from scratch on all historical data, while the cost of frequent retraining is unaffordable. In practice, we can perform rolling retraining once a month and adopt daily updates during the interval to learn new patterns timely. Our work is orthogonal to rolling retraining.
@@ B.2. Data Adaptation
Apart from model adaptation, data adaptation is another mainstream approach to concept drift in time series forecasting, which is orthogonal to model adaptation. The goal of data adaptation is to normalize historical training data and future test data into a common data distribution (RevIN; Dish-TS; SAN), reducing overfitting risks. Among them, RevIN (RevIN) is the most popular method that applies instance normalization to each time series lookback window and restores the statistics to the corresponding predictions, making each sample follow a similar distribution. Such normalization-based data adaptation methods mainly focus on the statistical changes in the mean and deviation of time series, while they overlook the distribution shifts in more complex temporal dependencies between time steps and spatial dependencies between variates (LIFT). Meanwhile, the forecast model may underfit the normalized time series, as the removed statistics can serve as informative signals for prediction (NSTransformer).
@@ C.1. Datasets
  • ETT (Informer) (Electricity Transformer Temperature) records seven features of electricity transformers from 2016 to 2018. ETTh1 is hourly collected and ETTm1 is 15-minutely collected.
  • ECL (DLinear) records the hourly electricity consumption of 321 clients from 2012 to 2014.
  • Weather (Informer) records 21 features of weather, e.g., air temperature and humidity in 2020.
  • Traffic (MTGNN) records the hourly road occupancy rates recorded by the sensors of San Francisco freeways from 2015 to 2016.
@@ C.2. Evaluation Settings
To the best of our knowledge, there is no open-source evaluation framework that considers the temporal gap in online learning for time series forecasting. Though Zhang et al. (OneNet) noticed the feedback delay, they sidestep this issue by performing one prediction task every $H$ time steps and performing the next task until the ground truth is available. Concretely, they use $\mathbf{X}_{t-L:t}$ to forecast $\mathbf{X}_{t+1:t+H}$ and then use $\mathbf{X}_{t-L+H:t+H}$ to forecast $\mathbf{X}_{t+H+1:t+2H}$. In contrast, the common practice is to perform forecasting at each time step between $t$ and $t+H$, continually updating the previous predictions of $\mathbf{X}_{t-L+H:t+H}$. The reason is that short-term forecasting is typically easier than long-term forecasting. Therefore, in this work, we implement a more realistic evaluation framework for online learning where we perform forecasting at each time step, which remains consistent with the traditional evaluation setting.
As for the ETTm1 dataset, popular works (e.g., Informer (Informer), PatchTST (PatchTST), and iTransformer (iTransformer)) usually use 69,680 time steps of the dataset. In contrast, FSNet and OneNet only used a quarter of it, i.e., 17,420 steps, and the training set only contains 3484 samples. This setting is in favor of small models (e.g., TCN) but is challenging for the state-of-the-art time series forecast models. Therefore, we take all 69,680 time steps following PatchTST and iTransformer. As such, the ETTm1 dataset in our experiment has the greatest number of samples, enriching the diversity of our datasets.
@@ C.3. Implementation Details
We conduct experiments on four Nvidia 4090 24GB GPUs. We ran each experiment 3 times with different random seeds and reported the average results. We follow the official implementation of the forecast models and the online learning methods, using their recommended model hyperparameters (e.g., the number of layers) and the Adam optimizer. In cases where default hyperparameter values are not provided, we conduct a grid search to find the optimal hyperparameters that yield the best performance. Following FSNet and OneNet, the lookback length $L$ for TCN is set to be $60$. For PatchTST and iTransformer, we search for the optimal lookback length and set $L$ to be $512$. We enhance TCN and FSNet by RevIN (RevIN) which is widely adopted by the state-of-the-art forecast models to reduce distribution shifts, lowering the forecasting loss in our experiments. As for the Optimal and Practical variants, we report the lowest errors when applying or not applying RevIN to FSNet and OneNet. Empirically, the Optimal variant without RevIN achieves better performance in some cases.
@@ E.2. Performance with Partial Ground Truth
[…]